### v1

# Robust certification of quantum instruments through a sequential communication game

Abstract
We propose a communication game in the sequential measurement scenario, involving a sender and two receivers with restricted collaboration among the latter parties. In the framework of the prepare-transform-measure scenario, we find a prominent quantum advantage in the receiver’s decoding of the message originally encoded by the sender. We show that an optimal trade-off between the success probabilities of the two receivers enables self-testing of the sender’s state preparation, the first receiver’s instruments, and the measurement device of the second receiver. Our protocol enables a more robust certification of the unsharp measurement parameter of the first receiver compared to the protocol without collaboration among the receivers. We further generalize our game to higher-dimensional systems, revealing greater quantum advantage with an increase in dimensions.
I Introduction
The study of the performance of quantum systems for communication and computational purposes gave birth to quantum information science [1, 2]. The study was essential as the technology was maturing towards manipulating single quantum systems [3]. In present times, the certification of quantum systems is regarded as an important cornerstone for the successful and secure implementation of quantum technology protocols. Robust self-testing forms the basis of certification of quantum states and measurements in a device-independent manner [4, 5, 6, 7, 8, 9, 10]. Practically relevant extensions in semi-device scenarios have also been proposed [11, 12, 13, 14, 15, 16]. The certification of quantum instruments plays a key role in the prepare-transform-measure scenario, forming an essential component of quantum communication in sequential networks.
Historically, considering the communication capacity of a photon as a carrier of classical information, Holevo derived a crucial result [17], which implies that a qubit cannot carry more than a bit of information. This apparent conclusion that a qubit is not more powerful than a bit was belied later due to a clever protocol that can reveal the power of a qubit over a bit. Specifically, the communication task of Random Access Codes (RAC) was developed on the conjugate coding protocol [18], in the prepare-and-measure scenario, which exhibits the supremacy of a qubit over a bit [19, 20, 21]. RAC is a fundamental protocol in information theory. In the case of classical $2\rightarrow 1$ RAC, two bits of classical information are compressed into a bit and are sent to a receiver, who has to retrieve any one of the bit values when asked randomly. The success probability of retrieving the encoded message is the figure of merit of the game. On the other hand, in the quantum version of random access code (QRAC), the encoding is done on a quantum system. It was shown that the success probability of correctly guessing a bit is larger when a qubit is a carrier than a bit in $2\rightarrow 1$ the RAC protocol [19, 20]. Later, quantum advantage has also been demonstrated for higher-dimensional QRAC [22].
QRAC has emerged as a basic building block in various quantum information processing tasks, such as semi-device-independent quantum key distribution [23], dimension witness [24, 25, 26], randomness generation [27, 28], quantum bidding [29], witnessing incompatibility of measurement [30]. The prepare and measure (PM) scenario enables certification of the quantum state and measurements [31, 32, 33, 34] in a semi-device-independent way, while device-independent self-testing [4, 5, 6, 7, 8, 9, 10] relies on the Bell test [35, 36]. In addition to various quantum information processing tasks, QRAC has been investigated in association with foundational issues of quantum mechanics as well [37, 38, 39, 40, 41]. Various experimental manifestations based on QRAC protocols have been proposed [37, 29, 42, 22, 43, 44, 45].
On the other hand, the idea of employing sequential measurements in investigating foundational and practical information processing issues have gained much importance in recent years [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. The sequential Bell test scenario has been employed to harness unbounded randomness from a single pair of entangled states [57] and self-testing quantum instruments as well [58]. Moreover, sequential measurements have been employed to study various other quantum correlations such as quantum steering [59, 60, 61], and sequential communication protocols have been suggested for information processing tasks such as entanglement detection, teleportation, and remote state preparation [62, 63, 64].
The generalization of the standard $2\rightarrow 1$ QRAC protocol to the sequential scenario has been performed [65, 66, 67], enabling various certification tasks. In [65], $2\rightarrow 1$ QRAC was extended to three parties where two independent receivers try to decode the bit position encoded earlier by the sender, randomly and sequentially. Here, the first decoder must have to measure weakly so that the second decoder can have a non-classical success probability. In this prepare-transform-measure (PTM) scenario, the success probability of the two decoders exhibits a trade-off, absent in the analogous classical counterpart. Such a protocol can be employed to self-test quantum instruments [65], and non-projective [68] and unsharp measurements [66] in a semi-device-independent manner. It may be noted that experimental implementations of sequential measurement protocols have also been performed [69, 70].
In the present work, we propose a new communication game enabling a more robust certification of quantum instruments. The standard $2\rightarrow 1$ QRAC scenario is extended to a tripartite communication game where there are two sequential receivers who are not completely independent. Here, the sender (Aparna) and the first receiver (Barun) act in the same way as in the standard RAC task, whereas the second receiver (Chhanda) has to guess the remaining bit position so that at the end of the game, both receivers together may have the full information about the message encoded by the sender. Specifically, in the context of $2\rightarrow 1$ QRAC, two bits of information are encoded into one bit by Aparna. Barun decodes the bit value of some position according to his random input. In our protocol, we allow restricted collaboration between two decoding receivers to make the task non-trivial and interesting as well. After decoding the asked bit value, Barun passes only his input information and the post-measured system to Chhanda without leaking any output information. We compute the optimal classical success probability of this proposed communication game and demonstrate an advantage through a quantum strategy.
We explore a trade-off between the success probability of Barun and Chhanda, which is absent in the classical case. The optimal trade-off leads to the certification of Barun’s quantum instrument together with Aparna’s state preparation and Chhanda’s measurement. We next investigate the robustness of the certification of quantum instruments in the presence of noise and estimate the lower and upper bounds on the sharpness parameter of Barun’s measurement. A significant outcome of our analysis is that our certification task is shown to be more robust compared to the approach of employing sequential QRAC without collaboration [65]. Since the gap between the lower and upper bounds of the estimated sharpness parameter of the first receiver’s measurement is narrower in our case, we are able to improve upon the earlier results in the context of our new communication game. We further generalize our game to the higher-dimensional $2^{d-1}\rightarrow 1$ QRAC scenario. We find the optimal classical success probability and numerically compute the quantum advantage up to dimension $6$.
The manuscript is organized in the following manner. We begin with the description of the new communication game in the next section. In Section III, we derive a trade-off relation between the success probabilities of the two receivers and show how this enables robust certification of the quantum components. Next, we generalize the proposed communication task to the higher-dimensional case in Section IV. We conclude with a summary of our analysis and results in Section V.
II Sequential game with restricted collaboration
Figure 1: Schematic diagram for a two-receiver communication game with restricted collaboration in ‘d’-dimensional quantum settings. Sender Aparna compressed a 2-dit message into a qudit and sends it to the first receiver, Barun, who decodes a dit position randomly. The task of the second receiver, Chhanda, is to guess the leftover information under the restriction that she can only use the knowledge of Barun’s input and has no information of his output from the communication.
In this section, we describe our proposed new communication game pertaining to the prepare-transform and measure scenario among three parties: Aparna, Barun, and Chhanda. Initially, two dits (higher-dimensional generalizations of a bit) of information are encoded into one dit by the sender, Aparna. The first receiver, Barun, tries to decode some partial message randomly from the encoded system and then passes it to Chhanda, who tries to retrieve the remaining message. The crucial point is that, unlike the protocol [65], here two receivers are not independent, and we allow some limited collaboration between them. In the next two subsections, we will discuss both the classical and quantum strategies to play this game, followed by results in the subsequent sections.
II.1 Two-receiver classical access code
In the classical version of this game, Aparna first selects a message randomly from $d^{2}$ options, which is a 2-dit string $x=(x_{0},x_{1})$. She encodes it into a one-dit message ‘m’ and sends it to Barun where, $x_{i},~{}m\in\{0,1,...,d-1\}$. Barun receives a random input $y\in\{0,1\}$ and needs to guess the $x_{y}$ dit value of Aparna’s message. After decoding the asked dit value, Barun will let Chhanda know about which position she has to decode and relay the post-measured system with a restriction. Here, the restriction is that the relayed system will carry no information about Barun’s output, i.e., Barun passes only his input information and the post-measured system to Chhanda without leaking any output information. Chhanda’s task is to infer the other dit, $x_{\bar{y}}$, of Aparna’s message (see, Fig.1).
In this communication task, Barun’s role is similar to that of the traditional RAC [21, 22]. To maximize Barun’s success probability of correctly retrieving the asked dit value, they can adopt the first-dit encoding strategy without loss of generality. With this approach, when Barun’s input is $y=0$, he retrieves the dit with probability 1. However, for the second dit (when $y=1$), his success probability is $\frac{1}{d}$ as he has no option other than a random guess. Therefore, on average, his success probability is $\frac{1}{2}\left(1+\frac{1}{d}\right)$. This classical strategy was conjectured to be optimal [22] and proved recently in [71].
Chhanda, however, lacks this advantage due to the restricted communication from Barun. As she only has the information about Barun’s input but not about his output, her success probability remains limited to $\frac{1}{d}$. The total classical success probability of retrieving the sender’s full encoded message, jointly by the two receivers, according to the above stated strategy of restricted collaboration, is given by,
$P^{C}(b=x_{y},c=x_{\bar{y}}|x,y)=\frac{1}{2d}\left(1+\frac{1}{d}\right).$ (1)
II.2 Two-receiver quantum access code
Now let’s describe the two receiver quantum access code with limited collaboration between the receivers. Aparna prepares a quantum state $\rho_{x}$ that encodes the selected classical message, $x=(x_{0},x_{1})$$\in\{0,1,...d-1\}^{2}$. Two receivers, Barun and Chhanda, only know the set of states of dimension $d$ from which Aparna selects in a single run of the game, but not the exact state. After preparing the state, Aparna sends it to Barun, who applies his measuring instrument, represented mathematically by the Kraus operators $\{K_{b|y}\}$. When this instrument acts on the state $\rho_{x}$, the outcome is $b\in\{0,1,\dots,d-1\}$, and the updated state according to the Lüdder transformation rule [72], is given by,
$\rho_{x}^{y,b}=\frac{K_{b|y}\rho_{x}K_{b|y}^{\dagger}}{\text{Tr}(\rho_{x}K_{b|y}K_{b|y}^{\dagger})}.$ (2)
To satisfy the completeness relation for Barun’s measurement, we require that $\sum_{b=0}^{d-1}M_{b|y}=\mathbb{I}$, where $M_{b|y}=K_{b|y}^{\dagger}K_{b|y}$ are the elements of a positive operator-valued measure (POVM). The correlation between Aparna and Barun is captured by the average success probability of correctly guessing the asked dit value, which can be obtained via Born’s rule in the following way. In the present work, we use the words average success probability, success probability, and figure of merit of the game interchangeably.
$P_{AB}^{Q}(b=x_{y}|x,y)$ $=\frac{1}{2d^{2}}\sum_{x,y}p(b=x_{y}|x,y)$ (3)
$=\frac{1}{2d^{2}}\sum_{x,y}Tr(\rho_{x}M_{x_{y}|y})$
According to the rules of the game, Barun will disclose his measurement choice, denoted by $y$, but not the corresponding outcome. Therefore, the post-measurement state received by Chhanda, is simply the average over all possible outcomes for a fixed input $y$, which is
$\tilde{\rho}_{x}^{y}=\sum_{b}K_{b|y}\rho_{x}K_{b|y}^{\dagger}.$ (4)
Now, to retrieve the leftover information, Chhanda will perform a generalized measurement $\{N_{0},N_{1}\}$ on $\tilde{\rho}_{x}^{y}$, where, $\sum_{b=0}^{d-1}N_{b|y}=\mathbb{I}$. Chhanda’s figure of merit is the success probability of correctly guessing the remaining part of the encoding and is given by the following way.
$P_{AC}^{Q}(c=x_{\bar{y}}|x,y)$ $=\frac{1}{2d^{2}}\sum_{x,\bar{y}}p(c=x_{\bar{y}}|x,\bar{y})$ (5)
$=\frac{1}{2d^{2}}\sum_{x,b,\bar{y}}\text{Tr}(K_{b|y}\rho_{x}K_{b|y}^{\dagger}N_{x_{\bar{y}}|\bar{y}}).$
The total average success probability of retrieving the full information of Aparna’s encoding by Barun and Chhanda together employing this restricted collaborative strategy is
$P^{Q}(b=x_{y},c=x_{\bar{y}}|x,y)=P_{AB}^{Q}P_{AC}^{Q}.$ (6)
III Quantum advantage and robust certification of instruments
In this section, we analyze the one sender, two receiver communication game in the presence of restricted collaboration between the receivers. We show explicitly that in quantum settings, an interesting non-trivial trade-off between the figure of merit of the two receivers arises, which leads to robust certification of quantum instruments. We consider the minimal scenario here, i.e., a two-bit message compressed into one bit, and generalize it to higher dimensions in the next section.
III.1 Trade-off between Barun’s and Chhanda’s respective figure of merit
Whatever strategy we follow, our goal is to maximize the success probability of the second receiver given that the first receiver obtains a fixed value of the non-classical success probability. This consideration characterizes the set of allowed regions in the correlation space. In the classical version of the game (Sec. II.1), Barun and Chhanda’s success probabilities are independent, resulting in a trivial scenario with no observable trade-off. In contrast, the quantum strategy (Sec. II.2) introduces a more complex interplay between these probabilities.
To formulate the problem in the quantum setting, we fix the value of $P_{AB}^{Q}$ (denoted as $\beta$) and seek to determine the maximum achievable value of $P_{AC}^{Q}$ over state preparations, instruments, and measurements under the relevant constraints by casting this problem as an SDP, which is explicitly given below. Our analysis focuses on the range $\beta\in\left[\frac{1}{2},\frac{1}{2}\left(1+\frac{1}{\sqrt{2}}\right)\right]$, which encapsulates the optimal trade-off between the success probability of two receivers. This range represents the nontrivial boundary of the quantum set of correlations in the $(P_{AB}^{Q},P_{AC}^{Q})$ plane. It is to be mentioned here that, unlike the sequential QRAC [65], here the second receiver has no random input, and together both Barun and Chhanda have a non-zero probability of decoding the full classical information encoded by Aparna.
Formally, this leads to the following optimization problem:
$P_{AC}^{Q,\beta}$ $=\max_{\rho,U,M,N}P_{AC}^{Q}$ (7)
s. t. $\forall x:\rho_{x}\in\mathbb{C}^{2},\;\rho_{x}\geq 0,\;\text{Tr}(\rho_{x})=1,$
$\forall z,c:N_{c|z}\geq 0,\;N_{0|z}+N_{1|z}=\mathbb{I},$
$\forall y,b:U_{yb}\in\mathrm{SU}(2),$
$\;M_{b|y}\geq 0,\;M_{0|y}+M_{1|y}=\mathbb{I},$
$\forall y,~{}~{}\sum_{x}\tilde{\rho}_{x}^{y}=\frac{\mathbb{I}}{2},$
and $\text{and }P_{AB}^{Q}=\beta.$
Figure 2: Trade-off between the success probabilities of two receivers is shown. The region on and below the red line is achievable employing a quantum strategy. Therefore, the red line refers to the optimal trade-off and hence, the quantum boundary. The deep dark line indicates the classical boundary, beyond which is the non-classical region.
In the above formulated SDP (7), we maximize Chhanda’s success probability by optimizing over all possible state preparations and measurements. The condition $P_{AB}^{Q}=\beta$ is enforced using suitable instruments, and its Kraus operator representation in the polar decomposition form is given as $K_{b|y}=U_{y|b}\sqrt{M_{b|y}}$, where $U_{y|b}$ represents a unitary operator, and $\{M_{b|y}\}$ is a POVM. The second last constraint is to impose the restriction that Chhanda has only the information about Barun’s input but no way to know his output.
As the game between Aparna and Barun is similar to the standard QRAC [21, 22], we can employ a quantum strategy for the above SDP (7), where Aparna encodes her preparations $\{\rho_{x}\}$ as,
$\rho_{00}$ $=\frac{1}{2}\left(\mathbb{I}+\frac{\sigma_{x}+\sigma_{z}}{\sqrt{2}}\right),\;\rho_{01}$ $=\frac{1}{2}\left(\mathbb{I}+\frac{\sigma_{x}-\sigma_{z}}{\sqrt{2}}\right),$ (8)
$\rho_{11}$ $=\frac{1}{2}\left(\mathbb{I}-\frac{\sigma_{x}+\sigma_{z}}{\sqrt{2}}\right),\;\rho_{10}$ $=\frac{1}{2}\left(\mathbb{I}-\frac{\sigma_{x}-\sigma_{z}}{\sqrt{2}}\right),$
and Barun performs an unsharp measurement, $M_{0}=\eta\sigma_{x}$, $M_{1}=\eta\sigma_{z}$ with sharpness parameter $\eta\in[0,1]$. Next, Chhanda performs $N_{0}=\sigma_{x}$, $N_{1}=\sigma_{z}$ measurements to optimize her success probability. Using this specified strategy, their success probability is given by,
$P_{AB}^{Q}$ $=\frac{1}{2}(1+\frac{\eta}{\sqrt{2}})$ (9)
$P_{AC}^{Q}$ $=\frac{1}{4}(2+\sqrt{2-2\eta^{2}})$
[…]
III.2 Self-testing based on optimal quantum strategy
[…]
An optimal pair of average quantum success probabilities of this multi-receiver quantum access protocol $(P_{AB}^{Q},P_{AC}^{Q})$ = $(\beta,P_{AC}^{\beta})$, as given by (9), self-tests the following physical entities
  • Aparna’s state preparations are as given in Eq.(8), which are pure and form a square on the Bloch sphere representation.
  • Barun’s measurement instruments are given by the Kraus operators $K_{b|y}=U_{y|b}\sqrt{M_{b|y}}$. The unitary freedom turns out to be the same for every observable and $M_{0}=\eta\sigma_{x},M_{1}=\eta\sigma_{z}$, where $\eta=\sqrt{2}(2\beta-1)$.
  • Chhanda’s rank-1 projective measurements $\{N_{0},N_{1}\}$ are given by $N_{0}=U\sigma_{x}U^{\dagger},N_{1}=U\sigma_{z}U^{\dagger}$.
III.3 Robust certification of sharpness parameter
In realistic conditions, all quantum systems deviate from the ideal target description, and hence, less than optimal success probabilities are observed, which can’t pinpoint the desired objects. Robustness analysis is the estimation of those objects for the suboptimal observed values. Therefore, any meaningful self-testing argument must come with a robustness analysis. Our above-mentioned self-testing statements in Sec.III.2 are valid only when the optimal trade-off point is achieved, which is not the case in all practical experimental conditions. Here we specifically consider the sharpness parameter of Barun’s measurement, which characterizes his instrument. We show how one can certify the sharpness parameter $\eta$, of Barun’s measurement in the presence of noise, in a more robust way than that in the approach based on sequential QRAC without collaboration [65]. We provide upper and lower bounds on the sharpness parameter for less than optimal observed success probabilities, which merge to the same value in noiseless conditions. Let’s state the result briefly here, and for the detailed analysis, we refer to the Appendix (B).
Figure 3: (a) Contour plots of $\delta$ and $\delta^{*}$ as functions of $\text{P}_{\text{AB}}^{Q}$ and $\text{P}_{\text{AC}}^{Q}$. Here, $\delta$ denotes the gap between the upper and lower bounds of the sharpness parameter estimated from the two-receiver communication game proposed in this work. On the other hand, $\delta^{*}$ is the bound gap drawn from the protocol without collaboration [65]. The noise configurations considered are : $p_{1}=0.95$, $p_{2}=0.93$, $p_{3}=0.95$; (b) Plots of $\delta$ and $\delta^{*}$ as functions of $\eta^{Target}$ denoting the value of the sharpness parameter to be self-tested. See also Table 1 and Appendix B.
[…]
$\eta\geq\sqrt{2}(2P_{AB}^{Q}-1).$ (10)
This bound is non-trivial whenever $P_{AB}^{Q}>1/2$. On the other hand, from the requirement that $P_{AC}^{Q}$ is non-classical, we can place an upper bound on $\eta$, given by
$\eta\leq\sqrt{8P_{AC}^{Q}-8(P_{AC}^{Q})^{2}-1}.$ (11)
Visibility
$(p_{1},p_{2},p_{3})$
Observed
Probabilities
$\delta$
$=\eta_{u}-\eta_{l}$
$\delta^{*}$
$=\eta_{u}^{*}-\eta_{l}^{*}$
(0.95, 0.9, 0.95)
$P_{AB}^{Q}=0.7138$
$P_{AC}^{*Q}=0.7826$
$P_{AC}^{Q}=0.7461$
0.1133 0.1964
(0.98, 0.95, 0.98)
$P_{AB}^{Q}=0.7328$
$P_{AC}^{*Q}=0.7955$
$P_{AC}^{Q}=0.7515$
0.0444 0.0824
(0.93, 0.88, 0.93)
$P_{AB}^{Q}=0.7046$
$P_{AC}^{*Q}=0.7726$
$P_{AC}^{Q}=0.7394$
0.1572 0.2617
Table 1: Comparison of estimation-error bands resulting from two different methods (with or without collaboration among the receivers) of self-testing the sharpness parameter. From Fig.3, we pick two different sets of values of noise in state preparation and measurements, and the obtained success probabilities of the corresponding games (second column). $P_{AB}^{Q}$ is the same in both protocols, as the task between the sender and the first receiver is the same as in the standard QRAC. $P_{AC}^{*Q}$ denotes the success probability of the second receiver, as obtained through the game without collaboration [65], whereas $P_{AC}^{Q}$ indicates our result. In the third column, $\delta$, the gap between the upper and lower bounds on the sharpness parameter resulting from our result is indicated, and the fourth column refers to the result obtained from the corresponding analysis in Ref. [65]. The greater robustness of self-testing of the sharpness parameter with respect to our proposed game with restricted collaboration between the receivers, compared to the standard sequential QRAC, is clearly revealed.
From the observed value of $(P_{AB}^{Q},P_{AC}^{Q})$, the pair, one can estimate the sharpness parameter, which lies within the upper and lower bounds. The narrower the gap between the two bounds, the better the certification. Specifically, we consider a noise model where Aparna’s preparation is given by $\rho_{x}^{{}^{\prime}}=p_{1}\rho_{x}+(1-p_{1})\frac{\mathbb{I}}{2}$, and the POVM elements corresponding to the measurement devices of Barun and Chhanda are represented as $M_{y}^{{}^{\prime}}=p_{2}M_{y}+(1-p_{2})\frac{\mathbb{I}}{2}$ and $N_{z}^{{}^{\prime}}=p_{3}N_{z}+(1-p_{3})\frac{\mathbb{I}}{2}$, respectively. Here, $p_{1}$ denotes the visibility of Aparna’s preparation device, $p_{2}$ and $p_{3}$ the visibilities of Barun’s and Chhanda’s measurement devices, respectively. In Fig. 3, we graphically illustrate the width of the plausible range for the sharpness parameter associated with the underlying noisy measurements.
In a noisy and suboptimal scenario, the sharpness parameter can be estimated within a significantly narrower range using our proposed communication game compared to the estimation based on the sequential QRAC protocol with no collaboration among the receivers [65]. For instance, suppose Aparna’s preparations have a visibility of 95%, while Barun’s and Chhanda’s measurements achieve visibilities of 93% and 95%, respectively. Under these conditions, our protocol estimates the sharpness parameter with an error gap of approximately $0.1133$, whereas the method in Ref. [65] leads to a larger error gap of $0.1964$. In Fig. 3(b), for the same visibility values, we plot the gap $(\delta,\delta^{*})$ as a function of $\eta^{\text{Target}}$, where $\eta^{\text{Target}}$ denotes the value of the sharpness parameter that an experimentalist may want to self-test. Detailed calculations and comparison plots for two other noisy scenarios are provided in Fig. 5 of Appendix B.
In Table 1, we present a comparison of the gap between the upper and lower bounds, estimated from the observed figures of merit in our communication game with restricted collaboration between the receivers and the protocol without such collaboration [65], for three different sets of noise levels. The gap between the bounds over the entire range of allowed observed success probabilities for the visibility set $(p_{1}=0.95,p_{2}=0.9,p_{3}=0.95)$ for the first set of noise levels is illustrated in Fig. 3(a). The results for the other two noisy scenarios are extracted from Fig. 5 in Appendix B. It can be seen that the protocol for self-testing the sharpness parameter based on restricted collaboration proposed by us is more robust compared to the one in Ref. [65] for a range of parameter values.
IV Sequential communication game in higher dimensions
Figure 4: Success probabilities as functions of the system dimension $d$. (a) Success probability for Barun. (b) Success probability for Channda. (c) Total success probability. The plots illustrate the quantum advantage of the proposed communication game. Notably, quantum strategies utilizing unsharp measurements outperform those restricted to sharp measurements. $\eta_{c}$ corresponds to the critical sharpness parameter.
Our proposed two-receiver communication game with restricted collaboration can be generalized to more sophisticated scenarios of QRAC with higher-dimensional systems and longer messages to be encoded or transmitted [21, 73]. Here we discuss the higher-dimensional generalization and compute numerically the lower bound of the optimal success probability employing the see-saw method up to dimension six (see Appendix C for the details). The see-saw is an iterative optimization method for affine functions of Hermitian operators, which was introduced initially to maximize quantum violation of a given Bell inequality [74].
The results are displayed in Fig. 4 where we separately show the success probalilities corresponding to Aparna, Barun and Chhanda, respectively, in the three sub-figures. In each of the sub-figures, we plot the respective success probabilities for three cases: (i) the classical success probability $P^{C}$, (ii) the quantum success probability at the threshold of quantum advantage $P(\eta=\eta_{c})$, with $\eta_{c}$ denoting the critical value at which Barun’s success probability matches the classical benchmark, given by $\eta_{c}=\frac{d-1}{d+\sqrt{d}-2}$, and (iii) the quantum success probability for sharp measurement by Barun, $P(\eta=1)$. The magnitudes of all the success probabilities are seen to drop with an increase in dimensions. Note from Figs. 4(a) and 4(b) that while Barun’s success probability drops with the increase of the sharpness parameter, the reverse occurs in the case of Chhanda’s success probability. Thus, the trade-off between Barun’s and Chhanda’s success is clearly maintained even in higher dimensions. From Fig. 4(c) we observe that the magnitude of quantum advantage increases with the dimension. Moreover, the resourcefulness of the unsharp measurement in the context of the present communication task is obvious, outperforming its sharp counterpart.
V Conclusions
Certification of quantum instruments [75] is a crucial ingredient of communication protocols relevant to quantum technology. Since noise is ubiquitous in most practical applications, the enhancement of robustness is an essential goal in such tasks. In the present work, we have proposed a new communication game extending the standard random access code paradigm to the sequential framework [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 59, 60, 61], involving one sender and two receivers with restricted communication allowed between the latter parties. Our protocol is based on sequential measurements in the prepare-transform-measure scenario, which has received much attention in recent times [31, 32, 33, 34, 65, 66, 67]. Through our proposed communication game, we can certify the sender’s state preparation and the measurement devices of the two sequential receivers.
It may be worth reemphasizing that quantum random access codes are fundamental communication tasks, which can exhibit quantum advantage even for the simplest quantum system—a qubit. Such advantage is one of the central issues in quantum information science, facilitating many quantum information processing tasks [23, 24, 25, 26, 27, 28, 29, 30], and is successful in addressing certain fundamental questions as well [37, 38, 39, 40, 41]. Non-trivial extensions of the standard QRAC enable self-testing of quantum instruments and non-projective measurements in the presence of noise in realistic experimental scenarios [43, 44, 45, 68]. Here we have proposed a novel three-party communication game based on $2\rightarrow 1$ QRAC, which involves collaboration among the receivers, and is hence, fundamentally different from the earlier extensions.
To summarize, in our proposed game, Aparna, the sender, encodes a two-bit message into a bit, and the task of the first receiver, Barun, is to decode a bit value according to his random input, whereas the second receiver, Chhanda, has to work out the leftover message. Thus, Barun and Chhanda may have some non-zero probability of correctly retrieving the full message, and they are allowed to have limited communication: Barun passes his input information to Chhanda, but there is no leakage of output information. Analyzing the protocol, we find a quantum advantage, which can certify Aparna’s state preparation, Barun’s instruments, as well as Chhanda’s measurement. Remarkably, the semi-device-independent certification of the sharpness parameter associated with Barun’s instrument is more robust in our protocol compared to that in the method based on sequential QRAC without any collaboration among the receivers [65]. We further generalize our sequential protocol to higher dimensions. By computing numerically the lower bound on the quantum success probability, we have shown that the quantum advantage increases with system dimensions, highlighting again the role of unsharp measurement in communication tasks relevant for sequential networks.
The analysis presented here gives rise to prospects for future work in several possible directions. It would be interesting to see how robust the certification task based on the proposed quantum access code under restricted communication is in higher dimensions. Further, it would be worthwhile investigating implications of our proposed game in other RAC-based tasks of foundational and information processing relevance [23, 24, 25, 26, 27, 28, 30, 37, 38, 39, 40, 41]. Moreover, our present analysis should motivate explorations of more sophisticated generalizations of the RAC paradigm in the sequential measurement framework, such as in multipartite scenarios involving more than two receivers, towards the goal of harnessing even greater quantum advantage with minimal resources.
Appendix A Optimal of success probability and self-testing
The success probability of the second receiver, $P_{AC}^{Q}$ is given by,
[…]
under the constraint that the success probability of the first receiver is,
[…]
We can use polar decomposition $K_{b|y}=U_{yb}\sqrt{M_{b|y}}$ and substitute $N_{1|z}=\mathbb{I}-N_{0|z}$, to get
[…]
Taking $\gamma_{y}=\sum_{x}(-1)^{x_{\bar{y}}}\rho_{x}$ and $D_{yb}^{\bar{y}}=U_{yb}^{\dagger}N_{x_{\bar{y}}|\bar{y}}U_{yb}$, we get
[…]
$=\frac{1}{2}+\max_{\rho,D,M}\frac{1}{8}\sum_{y,b}\lambda_{max}(\sqrt{M_{b|y}}\gamma_{z}\sqrt{M_{b|y}})$
and
$\gamma_{y}=\frac{1}{2}\left[(\eta_{00}-\eta_{11})+(-1)^{\bar{y}}(\eta_{01}-\eta_{10})\right]\cdot\vec{\sigma}$ (16)
Let $\vec{q}_{\bar{y}}=\frac{1}{2}\left[(\eta_{00}-\eta_{11})+(-1)^{\bar{y}}(\eta_{01}-\eta_{10})\right]$. Now,
$\forall M,\forall\vec{v}\in\mathbb{R}^{3}:\quad\sum_{b=0,1}\lambda_{max}[\sqrt{M_{b}}(\vec{v}\cdot\vec{\sigma})\sqrt{M_{b}}]\leq|\vec{v}|$ (17)
and
$P_{AC}^{Q}\leq\frac{1}{2}+\frac{1}{8}\sum_{y,b}\lambda_{\max}\left[\sqrt{M_{b|y}}(\vec{q}_{\bar{y}}\cdot\vec{\sigma})\sqrt{M_{b|y}}\right]$ (18)
Now define $T\equiv\sum_{y,b}\lambda_{\max}\left[\sqrt{M_{b|y}}(\vec{q}_{\bar{y}}\cdot\vec{\sigma})\sqrt{M_{b|y}}\right]$. Let’s consider an observable $M_{y}=M_{0|y}-M_{1|y}$. We can write $M_{y}=g_{y0}\mathbb{I}+\vec{g}_{y}\cdot\vec{\sigma}$ where $\vec{g}_{y}=(g_{y1},g_{y2},g_{y3})$ with $|\vec{g}_{y}|\leq 1$ and $|\vec{g}_{y}|-1\leq g_{y0}\leq 1-|\vec{g}_{y}|$. This constraint ensures the operator is positive:
$M_{b|y}=f_{yb}\ket{\vec{g}_{y}}\bra{\vec{g}_{y}}+h_{yb}\ket{-\vec{g}_{y}}\bra{-\vec{g}_{y}}$ | (19)
Here, |$\ket{\vec{g}_{y}}$ is the pure state corresponding to the bloch sphere direction $\vec{g}_{y}$,
$f_{yb}=\frac{1}{2}(1+(-1)^{b}g_{y0}+(-1)^{b}|\vec{g_{y}}|)$ (20)
$h_{yb}=\frac{1}{2}(1+(-1)^{b}g_{y0}-(-1)^{b}|\vec{g_{y}}|)$
Now, we can write Eq. (13) as,
$P_{AB}^{Q}=\beta=\frac{1}{8}(4+|\vec{q_{0}}|g_{01}+|\vec{q_{1}}|g_{13})$ (21)
Then, solving the characteristic equation $\text{det}\left(\sqrt{M_{b|y}}(\vec{q}_{\bar{y}}\cdot\vec{\sigma})\sqrt{M_{b|y}}-\nu\mathbb{I}\right)=0$, we get,
$T=\sum_{y,b}\frac{|\vec{q}_{\bar{y}}|}{2}\sqrt{(1+(-1)^{b}g_{y0})^{2}-|\vec{g}_{y}|^{2}(1-\bra{\vec{g}_{y}}\hat{q}_{\bar{y}}\cdot\vec{\sigma}\ket{\vec{g}_{y}}^{2})}$ (22)
We now perform the optimization over $g_{y0}$ by considering the two terms corresponding to $y=0,1$ separately. This amounts maximizing an expression of the form $\sqrt{(1+w)^{2}-z}+\sqrt{(1-w)^{2}-z}$, for positive $z$. It can be shown such that such a function is uniquely maximized by setting $w=0$. So we require $g_{00}=g_{10}0$. Moreover, $(\vec{q}_{0},\vec{q}_{1})$ has no component along y-axis, and one can optimally choose $g_{02}=g_{12}=0$. This enables us to express
$\max T=|\vec{q}_{0}|\sqrt{1-(g_{11}^{2}+g_{13}^{2})(1-g_{11}^{2})}+|\vec{q}_{1}|\sqrt{1-(g_{01}^{2}+g_{03}^{2})(1-g_{03}^{2})}$ (23)
Note that $g_{03}$ and $g_{11}$ do not appear in the constraint, and that they are associated with different settings of Barun, and appear in different terms in Eq. (23). Thus we can separately maximize using standard differentiation. This returns the unique maximum attained for $g_{03}=g_{11}=0$, and hence, we have,
$P_{AC}^{Q}\leq\frac{1}{2}+\frac{1}{8}(|\vec{q}_{0}|\sqrt{1-g_{13}^{2}}+|\vec{q}_{1}|\sqrt{1-g_{01}^{2}})\equiv\tilde{P}$ (24)
Let $g_{01}=\cos\zeta_{0}$ and $g_{13}=\cos\zeta_{1}$ for $\zeta_{0},\zeta_{1}\in[0,\frac{\pi}{2}]$, so that
$\tilde{P}=\frac{1}{2}+\frac{1}{4}(\cos\frac{\theta}{2}\sin\zeta_{1}+\sin\frac{\theta}{2}\sin\zeta_{0})$ (25)
By putting the relevant values we can rewrite $\beta$ as,
$\beta=\frac{1}{8}(4+\cos\frac{\theta}{2}\cos\zeta_{0}+\sin\frac{\theta}{2}\cos\zeta_{1})$ (26)
Lemma 1: For every tuple $(\theta,\zeta_{0},\zeta_{1})$ corresponding to $(\beta,\tilde{P})$, there exists another tuple $(\theta,\zeta_{0},\zeta_{1})=(\frac{\pi}{2},\zeta,\zeta)$ that produces $(\beta,\tilde{P}^{{}^{\prime}})$ with $\tilde{P}^{{}^{\prime}}\geq\tilde{P}$, Moreover, $\theta=\frac{\pi}{2}$ and $\zeta_{0}=\zeta_{1}$ leads to the optimal $\tilde{P}^{{}^{\prime}}$.
Proof:
To prove this lemma, we must show that $\forall\theta,\zeta_{0},\zeta_{1}\in[0,\frac{\pi}{2}]$ there exists $\zeta\in[0,\frac{\pi}{2}]$ such that,
$\cos\frac{\theta}{2}\cos\zeta_{0}+\sin\frac{\theta}{2}\cos\zeta_{1}=\sqrt{2}\cos\zeta$ (27)
$\cos\frac{\theta}{2}\sin\zeta_{0}+\sin\frac{\theta}{2}\sin\zeta_{1}\leq\sqrt{2}\sin\zeta$
By squaring both equations and inequalities of Eq. (27), we can combine them to eliminate $\zeta$:
$\cos\theta(\cos^{2}\zeta_{0}-\cos^{2}\zeta_{1})+\sin\theta\cos(\zeta_{0}-\zeta_{1})\leq 1$ (28)
Differentiating w.r.t. $\zeta_{0}$, one finds that the optimum of the left-hand side is attained for $\zeta_{1}=\zeta_{0}$, which proves the above relation.
From the lemma 1, we set $\theta=\frac{\pi}{2}$ in (25) and (26), and $g_{01}=g_{13}=g$. Then,
$g=\sqrt{2}(2\beta-1).$ (29)
We also have $\tilde{P}=\frac{1}{2}+\frac{\sqrt{1-g^{2}}}{2\sqrt{2}}$, and hence,
$P_{AC}^{Q,\beta}\leq\frac{1}{4}(2+\sqrt{16\beta-16\beta^{2}-2}).$ (30)
The r.h.s. of Eq.(30) corresponds to the success probabilities emanating from the quantum strategy (9) in the main text, showing that the above upper bound could be saturated with an explicit quantum strategy.
In order to analyze the self-testing argument explicitly, let us more closely examine the steps needed to arrive at Eq.  (30). First, we have already shown that the preparations must be pure, pairwise antipodal and by lemma 1, they must have a relative angle of $\frac{\pi}{2}$. This corresponds to a square in a disk of the Bloch sphere. The above arguments fully characterize Aparna’s preparations up to a reference frame. For Barun’s instrument, we have shown that the Bloch vectors ($\vec{g}_{0},\vec{g}_{1}$) can only have non-zero components in the $x$- and $z$-directions respectively, and that the length of the Bloch vector is given by Eq. (29). This fully characterizes the Bloch vectors. Moreover, in Eq. (15), we require that $D_{yb}^{\bar{y}}$ is aligned with the eigenvector of $\sqrt{M_{b|y}}D_{yb}\sqrt{M_{b|y}}$ corresponding to the largest eigenvalue. In case of Chhanda’s measurement, we now have $\gamma_{0}=\sigma_{x}$ and $\gamma_{1}=\sigma_{z}$, while $M_{0}\propto\sigma_{x}$ and $M_{1}\propto\sigma_{z}$. Therefore, it follows that $\forall y,b:D_{yb}^{0}=|+\rangle\langle+|$ and $\forall y,b:D_{yb}^{1}=|0\rangle\langle 0|$, we have,
$U_{yb}^{\dagger}N_{0|0}U_{yb}=\ket{+}\bra{+}$ | (31)
$U_{yb}^{\dagger}N_{0|1}U_{yb}=\ket{0}\bra{0}$ |
This implies that all unitaries are equal: $U_{yb}=U$. Therefore, Chhanda’s observables $C_{z}=C_{0|z}-C_{1|z}$ satisfy $C_{0}=U\sigma_{x}U^{\dagger}\quad\text{and}\quad C_{1}=U\sigma_{z}U^{\dagger}.$ This completes our self-testing argument
Appendix B Upper and lower bound on the sharpness parameter $\eta$
Figure 5: (a) Contour plots of $\delta$ and $\delta^{*}$ as functions of $\text{P}_{\text{AB}}^{Q}$ and $\text{P}_{\text{AC}}^{Q}$ for $p_{1}=0.98$, $p_{2}=0.95$, $p_{3}=0.98$. (b) Plots of $\delta$ and $\delta^{*}$ as functions of $\eta^{\text{Target}}$ for the same visibility values. (c) Contour plots of $\delta$ and $\delta^{*}$ as functions of $\text{P}_{\text{AB}}^{Q}$ and $\text{P}_{\text{AC}}^{Q}$ for $p_{1}=0.93$, $p_{2}=0.88$, $p_{3}=0.93$. (d) Plots of $\delta$ and $\delta^{*}$ as functions of $\eta^{\text{Target}}$ for the same visibility values. Here, $\delta$ denotes the gap between the upper and lower bounds of the sharpness parameter estimated from the multi-receiver communication game proposed in this work, while $\delta^{*}$ represents the corresponding bound-gap reported in Ref. [65].
In order to bound the sharpness of Barun’s instrument, consider Barun’s success probability, which using Eq. (21), is given by,
$P_{AB}^{Q}=\frac{1}{8}(4+|\vec{g}_{0}||\vec{q}_{0}|\hat{q}_{0}\cdot\hat{g}_{0}+|\vec{g}_{1}||\vec{q}_{1}|\hat{q}_{1}\cdot\hat{g}_{1})$ (32)
For simplicity, we can now fix the sharpness parameter of Barun’s two settings as, $\eta\equiv|\vec{g}_{0}|=|\vec{g}_{1}|$. It follows that,
$\eta=\frac{8P_{AB}^{Q}-4}{|\vec{g}_{0}|\hat{q}_{0}\cdot\hat{g}_{0}+|\vec{g}_{1}|\hat{q}_{1}\cdot\hat{g}_{1}}$ (33)
The lower bound of $\eta$ is obtained by maximizing the denominator. This occurs when $\hat{q}_{0}\cdot\hat{g}_{0}=\hat{q}_{1}\cdot\hat{g}_{1}=1$ and $|\hat{q}_{0}|=|\hat{q}_{1}|$, and one gets
$\eta\geq\sqrt{2}(2P_{AB}^{Q}-1)$ (34)
For obtaining the upper bound of $\eta$, we make use of the optimal choice of the success probability from Eq.9, given
$P_{AC}^{Q}\leq\frac{1}{2}+\frac{\sqrt{1-\eta^{2}}}{2\sqrt{2}}$ (35)
Inverting the above inequality, leads to,
$\eta\leq\sqrt{8P_{AC}^{Q}-8(P_{AC}^{Q})^{2}-1}$ (36)
Appendix C Success probabilities for higher dimensional systems
Here we employ the technique of the see-saw algorithm, which is especially well-suited to our optimization problem required to generalize the analysis in Section II.2 to higher dimensional systems. Our goal here is to maximize the success probabilities over $d$-dimensional quantum states $\rho_{x}$ and $d$- outcome POVMs. The success probability of the first receiver is obtained as,
$\text{max}\;P_{AB}^{Q}=\frac{1}{2d^{2}}\sum_{x,y}Tr(\rho_{x}M_{b=x_{y}|y})$ (37)
subject to $\forall x:\rho_{x}\in\mathbb{C}^{d},\;\rho_{x}\geq 0,\;\text{Tr}(\rho_{x})=1,$
$\forall y,b:\;M_{b|y}\geq 0,\;\sum_{b=0}^{d-1}M_{b|y}=\mathbb{I}.$
The objective function in the above maximization problem (37), is a product of semi-definite matrices. The see-saw algorithm is an iterative optimization technique that alternates between fixing one matrix and optimizing over the other. For the sake of completeness let’s describe the algorithm briefly by considering the maximization of $P_{AB}^{Q}$. First, we randomly select an initial state for Aparna’s preparation and proceed to optimize the objective function over Barun’s POVMs.
$\text{max}\;P_{AB}^{Q}=\frac{1}{2d^{2}}\sum_{x,y}Tr(\boldsymbol{\rho_{x}}M_{b=x_{y}|y})$ (38)
s. t. $\forall y,b:\;M_{b|y}\geq 0,\;\sum_{b=0}^{d-1}M_{b|y}=\mathbb{I}.$
Symbols in bold indicates which variable is kept fixed. Since the objective function is linear in the semi-definite variables, the problem is an SDP. In the second step we fix the optimal measurements for Barun, obtained from the previous SDP, and optimize over Aparna’s preparations, i.e.,
$\text{max}\;P_{AB}^{Q}=\frac{1}{2d^{2}}\sum_{x,y}Tr(\rho_{x}\boldsymbol{M_{b=x_{y}|y}})$ (39)
s. t. $\forall x:\rho_{x}\in\mathbb{C}^{d},\;\rho_{x}\geq 0,\;\text{Tr}(\rho_{x})=1.$
This problem is again an SDP. Now in the first step of the next iteration, Aparna’s optimal state preparation, coming out of the last optimization problem, is kept fixed and the SDP (38) is run followed by (39). In this way the algorithm is run in a loop several times until it converges to an apriori fixed value of precision. While this approach often yields good results, converge to a global maximum is not guaranteed. To increase the likelihood of finding a global maximum, the entire procedure is repeated multiple times with different randomly chosen initial states.
We can apply the above stated see-saw technique to optimize the total success probability of the proposed communication game. Here, Barun performs an unsharp measurement of the form $\tilde{M}_{b|y}=\eta M_{b|y}+(1-\eta)\frac{\mathbb{I}}{d}$ with sharpness parameter $\eta\in[0,1]$. A natural choice for the Kraus operator leads to Lüders’ instrument [72], which was proven to be optimal in the context of measurement-information gain trade-off [47], is given as,
$K_{b|y}=\left(\sqrt{\frac{1+(d-1)\eta}{d}}-\sqrt{\frac{1-\eta}{d}}\right)M_{b|y}+\sqrt{\frac{1-\eta}{d}}\mathbb{I}.$ (40)
For this unsharp measurement, for a given value of $\eta$, the success probability of Barun and subsequently that of Chhanda can be computed using the see-saw method to obtain a lower bound on the optimal value.
Let’s describe the relevant steps employing the see-saw algorithm to compute the total success probability of our two-receiver communication game with restricted collaboration.
Step 1: We initialize Aparna’s state preparation, $\{\boldsymbol{\rho_{x}}\}$, with random quantum states and proceed to maximize the objective function for a given value of $\eta$ over Barun’s measurements.
$\text{max}\;P_{AB}^{Q}=\frac{1}{2d^{2}}\sum_{x,y}Tr(\boldsymbol{\rho_{x}}\tilde{M}_{b=x_{y}|y})$ (41)
subject to $\forall y,b:\;\tilde{M}_{b|y}\geq 0,\;\sum_{b=0}^{d-1}\tilde{M}_{b|y}=\mathbb{I}.$
Step 2: Then apply the Kraus operators as Eq. (40) on the randomly chosen states in the previous step and maximize Chhanda’s success over her POVM.
$\text{max}\;P_{AC}^{Q}=\frac{1}{2d^{2}}\sum_{x,b,\bar{y}}\text{Tr}(K_{b|y}\boldsymbol{\rho_{x}}K_{b|y}^{\dagger}N_{x_{\bar{y}}|\bar{y}})$ (42)
subject to $\forall y,b:\;N_{b|\bar{y}}\geq 0,\;\sum_{b=0}^{d-1}N_{b|\bar{y}}=\mathbb{I}.$
Step 3: We fix the optimal POVMS obtained from step 1 as Barun’s measurements ($\boldsymbol{\tilde{M}_{b=x_{y}|y}}$) and optimize Aparna’s preparations to maximize the objective function.
$\text{max}\;P_{AB}^{Q}=\frac{1}{2d^{2}}\sum_{x,y}Tr(\rho_{x}\boldsymbol{\tilde{M}_{b=x_{y}|y}})$ (43)
subject to $\forall x:\rho_{x}\in\mathbb{C}^{d},\;\rho_{x}\geq 0,\;\text{Tr}(\rho_{x})=1.$
These three steps are repeated iteratively until convergence in both $P_{AB}^{Q}$ and $P_{AC}^{Q}$ is achieved.